OpenStudy (anonymous):

sound levels in decibels (db) can be computed by f(x) = 160+10 log x, where x is the intensity of the sound in watts per sq centimeter. The intensity of a sound is 10^-8 watts per square centimeter. How many decibels does this quantity represent. What is the sound level

OpenStudy (anonymous):

f(10^-8) = 160 + 10 * log(10^-8) =160 + 10 * -8 =160 - 80 = 80 dB.